OpenStudy (meilendurcer):

For the function y = f(x), what is the ordered pair for the point on the graph when x = p + 1? (x, f(p)) (x, p + 1) (p + 1, f(p + 1)) (p + 1, f(p))

OpenStudy (meilendurcer):

B?

ganeshie8 (ganeshie8):

Ordered pair : \((x, y)\)

ganeshie8 (ganeshie8):

since y = f(x), it wud be : \((x, f(x))\)

ganeshie8 (ganeshie8):

plugin \(x = p+1\) above ^

OpenStudy (meilendurcer):

so the first one

OpenStudy (meilendurcer):

not positive on that though

ganeshie8 (ganeshie8):

not quite, when u plugin \(x = p+1\) you get : \((p+1, f(p+1))\)

OpenStudy (meilendurcer):

ohhhh

ganeshie8 (ganeshie8):

just notice that we always put the "x" value first and "y" value second

OpenStudy (meilendurcer):

right

OpenStudy (meilendurcer):

this one also? Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one halfx + 5. y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4